Use the format to solve:

def count_consecutive(input_string):
output_list = []
#Write your code here
return output_list

#You may modify the code below for testing
input_string='occurred'
output_list = count_consecutive(input_string)
print(output_list)

Write a Python program which takes an input_string as input parameter and returns an output_list as per the logic below –

  • For each character in the input_string
  • If the character is occurring consecutively,
    • Count the number of consecutive occurrences of the character
    • Concatenate the character and the count of the consecutive occurrences of the character in input_string and add the result to the ouput_list

Example 1: If the input_string is "occurred", the count of consecutive occurrence of characters 'c' and 'r' are 2. Hence the output_list would be ['c2', 'r2']

Example 2: If the input_string is "possessive", the count of consecutive occurrence of character 's' occurring at positions 2 and 5 are 2. Hence the output_list would be ['s2', 's2']

  • If there is no consecutive occurrence of any character in the given input_string add "X" (upper case) to the output_list

Example: In the input_string "Madam", there is no character occurring consecutively. So the output_list would be ["X"]

  • Return output_list

Assumptions:

  • The input_string would have at least one character
  • All the characters in the input_string would be alphabets in lowercase

Note: No need to validate assumptions

Sample Input and Output
Answer #1

def count_consecutive(input_string):
   output_list = []
   if(len(input_string)==1):
       output_list.append("X")
       return output_list
   i=0
   while(i<len(input_string)):
       j=i
       count=0
       s=""
       while(j<len(input_string)):
           if(input_string[i]==input_string[j]):
               count=count+1
               s=input_string[i]
               j=j+1
           else:
               break          
       if(count>1):
           s=s+str(count)
           output_list.append(s)
       i=i+1
   if(output_list==[]):
       output_list.append("X")
   return output_list

print(count_consecutive("ouccurred"))
print(count_consecutive("possessive"))
print(count_consecutive("p"))
print(count_consecutive("posesve"))